B. Find the finite population correction factor given the following: 1. ๐‘=150 ๐‘›=15 2. ๐‘=600 ๐‘›=35ย 


1
Expert's answer
2022-06-02T13:51:45-0400

1.


"\sqrt{\dfrac{N-n}{N-1}}=\sqrt{\dfrac{150-15}{150-1}}\approx0.9519"

2.


"\sqrt{\dfrac{N-n}{N-1}}=\sqrt{\dfrac{600-35}{600-1}}\approx0.9712"